OpenStudy (anonymous):

Use the Law of Sines to solve for all possible triangles that satisfy the given conditions. (If an answer does not exist, enter DNE. Round your answers to one decimal place. Below, enter your answers so that ∠A1 is smaller than ∠A2.)

OpenStudy (anonymous):

how do you know when theres a second triangle possible ??

OpenStudy (anonymous):

Where is the question?

OpenStudy (anonymous):

b = 29, c = 38, ∠B = 25°

OpenStudy (anonymous):

It says :satisfy the given condition, But no condition is given.

OpenStudy (anonymous):

i just put it : b = 29, c = 38, ∠B = 25°

OpenStudy (anonymous):

Do you know the law of sines?

OpenStudy (anonymous):

yes : sin A/a=sinB/b=sinC/c

OpenStudy (anonymous):

So have you solved for angles C and A and side a?

OpenStudy (anonymous):

yes i got C=146.4 and A=8.6 a=10.3

OpenStudy (anonymous):

so when it comes down to have this question on the test how would i know that i have to solve for a second triangle?

OpenStudy (anonymous):

h<a<b : two triangles where h = height.

OpenStudy (anonymous):

a >= b: one triangle a=h one triangle

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):



OpenStudy (anonymous):

You only need to solve for one triangle on your problem.

OpenStudy (anonymous):

okay i get it

OpenStudy (anonymous):

with the calculations from the first triangle how do i know if a second triangle is possible

OpenStudy (anonymous):

Can the height of your triangle be greater than a or b. Not if you have an obtuse triangle. Draw the triangle out and see.